# Unsupervised and Source-Free Ranking of Biomedical Segmentation Models

### v2

@@ 5.1 Semantic Segmentation
Tab. 1 shows all CTE variants exhibit strong linear and monotonic correlations with performance across all eight target datasets (four mitochondrial and four nuclei). Comparable results between input- and feature-space perturbations confirm both approaches—previously explored in uncertainty estimation [43, 39], model generalisation [15, 2, 61], and semi-supervised learning [48]—provide effective transferability estimates reflecting the robustness and boundary proximity of target embeddings. For feature-space perturbations, we examined where in the network they should be applied. Prior work [43, 25] suggests intermediate layers, where information is most redundant. To test generality, we applied spatial dropout either to all layers (Mito) or only to the bottleneck (Nuclei). As shown in Tab. 1, both strategies yielded strong correlations. Finally, since our model sets include networks trained both with and without augmentations, the consistently high CTE performance indicates that CTE is robust to training schemes and unaffected by source-domain augmentations during test-time perturbation.
In contrast to the classification case [15], “hard” (CTE-NHD) and “soft” consistency (CTE-EI) metrics perform equivalently well for semantic segmentation.
Uncertainty estimation methods [43] typically quantify output variance across multiple perturbation runs. For transferability ranking, we found that measuring consistency under a single perturbation per test image was sufficient, enabling lightweight model selection with only two inference passes per image. For both input and feature-space perturbations, a broad range of perturbation strengths effectively predicted transfer success via CTE (see Supp.), so all tables show results for single perturbation strengths.
Fig. 3 illustrates CTE performance on a representative target dataset (EPFL mitochondria segmentation). The results show a strong correlation across a diverse set of source models comprising three architectures (U-Net, Residual U-Net, and UNETR), trained on four source datasets (EPFL, Hmito, Rmito, and VNC) with and without input augmentations. All models perform similarly well on their respective source domains, yet visual similarity between source and target images does not reliably predict transferability. Transfer performance is clearly influenced by both architecture and training setup, as seen in the contrast between UNETR (2) and U-Net (3), trained on the same Rmito source data. Our consistency-based approach, by relying on frozen model weights, inherently accounts for both model characteristics and source-domain distributions, enabling a direct assessment of their suitability for a given target dataset.
While CTE generally correlates strongly with transfer performance, pathological cases can cause model mis-rankings. These occur primarily among low-performing transfers, where models exhibit “mode collapse”, producing overly uniform segmentations, such as all background or all foreground. Such degenerate outputs yield artificially high consistency. This limitation could be reduced by pairing consistency with prediction diversity or entropy, which helps distinguish truly stable models from those with collapsed outputs.
@@
When it comes to highly multi-class 3D settings (Tab. 2), both our CTE-EI and CTE-NHD metrics remain reliable, although CTE-NHD is slightly better, probably due to unreliable model calibration in a challenge setting [50]. Correlation scores are computed after aggregating semantically similar classes, i.e., averaging left/right Inferior Alveolar Canals into IACs, the 32 teeth into Teeth, and left/right Maxillary Sinuses into Sinus. We also provide an Overall score, which is the mean across all 42 classes in the ToothFairy2 dataset. Once again, both input and feature space perturbations yield strong and coherent rank correlations. Splitting the class-grouped results reveals our metrics are effective and stable across a morphologically heterogeneous set of classes.
| Metric | Cells | Nuclei | |||||
| K | S | P | K | S | P | ||
| CTE-ARS (Gauss) | 0.69 | 0.83 | 0.90 | 0.72 | 0.83 | 0.79 | |
| ±0.15 | ±0.09 | ±0.04 | ±0.25 | ±0.17 | ±0.21 | ||
| CTE-ARS (DropOut) | 0.69 | 0.83 | 0.91 | 0.80 | 0.85 | 0.84 | |
| ±0.09 | ±0.06 | ±0.04 | ±0.28 | ±0.24 | ±0.11 | ||
Unsupervised Validation. UDA can improve model transfer performance, but if multiple models are adapted, the user again faces the problem of ranking the (fine-tuned) models. Tab. 3 shows the performance of CTE on this task, compared to state-of-the-art Transfer Score. We use mitochondria datasets and two simple UDA methods: a classic Mean Teacher (MT)-based self-supervised domain adaptation with source model predictions as pseudo-labels [65] and adapted batch normalisation (AdaBN) [35]. We observe overall high performance with strong correlation between CTE and final UDA model performance over three out of four target datasets for both UDA approaches. CTE outperforms the previous state-of-the-art for UDA validation, Transfer Score (TS) [78]. The VNC dataset exhibits weaker correlations for both CTE and TS metrics, particularly under the Mean Teacher fine-tuning approach. For CTE, correlation scores are adversely affected by a few outlier models, notably those trained on the EPFL source data. Further inspection revealed these to be pathological transfer cases, where the model consistently segments non-mitochondrial structures in the target data, yielding stable but incorrect predictions and, consequently, low task performance. This misalignment likely arises from distributional differences between the source and target datasets and from the inherent visual similarity of distinct structures in microscopy images, which can induce class confusion during transfer. Consistency regularisation in Mean Teacher training further reinforces this erroneous behaviour. In these cases, the implicit assumption that source and target models address the same task is violated, so models can be ranked well under direct transfer but exhibit outlier behaviour after self-training. The ranking scenario after UDA is inherently more challenging, as the target data effectively becomes part of the model’s adapted source distribution, reducing the impact of perturbations used for CTE estimation. Moreover, the VNC dataset is the smallest of all evaluated targets, containing only 20 slices in total, which further limits statistical robustness and contributes to the weaker correlations observed. As shown in Tab. 3, the AdaBN UDA approach—lacking a consistency component—remains less affected, resulting in higher correlation scores.
@@ 5.2 Instance Segmentation
To our knowledge, we are the first to address this problem, so no baseline is available. Unlike the semantic case, instance segmentation does not produce a fixed set of masks with direct correspondence to feature vectors, hence previously compared transferability estimators cannot be applied. For self-trained models we follow a popular instantiation process [53] of predicting cell boundaries and then using seeded watershed and graph partitioning algorithm GASP [6] to obtain instances. Tab. 4 shows the average CTE-ARS performance for instance segmentation, across two tasks and seven target datasets, where it achieves a strong linear and monotonic correlation for all targets.
Evaluating instance model transfer through output prediction consistency enables direct comparison across instance segmentation models employing different instantiation strategies. Fig. 4 presents the transfer ranking of several publicly available models on the Covid-IF cell segmentation dataset: [3] with three transformer backbones (ViT-T, ViT-B, ViT-L), Cellpose-SAM [51], and “Powerful-Chipmunk” [53], a specialist U-Net model trained on this dataset and released via the BioImage Model Zoo [49]. These models differ substantially in architecture and instance formation methodology. The Covid_IF dataset was not included in the training data of , while inclusion for Cellpose-SAM is uncertain. CTE accurately reproduces the true performance ranking across this diverse model set (see Fig. 4).111The ranking does not show the superiority of any model, but is a target dataset specific ranking of automatic segmentation performance.
@@ 6 Limitations and Conclusion
The proposed consistency metric assumes source-target task alignment. For generalist models like and Cellpose-SAM, this may not fully hold. The models are finetuned for both cell and nuclei and have no automated mechanism to control which structure is segmented. While supports prompt-based guidance, it requires human input and thus falls outside the scope of automated model ranking. In datasets like Covid_IF, a channel can implicitly contain information about both classes. Models occasionally switch to segmenting nuclei instead of cells (see Supp.). This reduces cell segmentation performance but not consistency scores, as predictions remain internally stable.
Furthermore, the approach can only correlate with pixel-wise metrics that treat all pixels equally, such as Dice or Adapted Rand Error. It is less aligned with boundary-sensitive metrics that assign greater importance to specific regions. In future work, the method could be extended to incorporate pixel-importance weighting within the consistency calculation to address this limitation.
In conclusion, we introduced a novel zero-shot transferability estimator for segmentation models that requires no source training data or target domain labels. Our estimator rankings strongly correlate with direct model performance rankings, achieving state-of-the-art results for semantic and instance segmentation transferability estimation. Our estimators require only two inference runs on test data and apply to blackbox and greybox models, facilitating efficient model selection.
7 Acknowledgements
The research was carried out as part of the AI4Life consortium. AI4Life receives funding from the European Union’s Horizon Europe research and innovation programme under grant agreement number 101057970. Special thanks also to Fynn Beuttenmueller for many constructive conversations.
@@ 7 Acknowledgements
8 Datasets
The general experimental setup for source model transferability ranking is as follows, a set of ‘source’ models are trained for the same task (i.e nuclei semantic segmentation). The models can have a range of architectures and source datasets. A single dataset is then selected as ‘target’ and all of the models from are applied to the test set of the target data. We then rank the transferability of the models and compare this ranking to the transfer performance ranking, obtained using ground-truth. In this section we further detail the employed datasets, and in section Sec. 9 the tested models.
We experimented applying our consistency based transferability estimator (CTE) across a wide range of datasets and modalities from the microscopy and biomedical domains. We focused on ranking the transferability of semantic and instance segmentation models across several distinct tasks. The tasks covered arguably the most common problems in microscopy image analysis; nuclei, cell and mitochondria segmentation. We additionally investigated a popular biomedical segmentation challenge (i.e, ToothFairy2) for human jaw segmentation. Our wide range of experiments across such a diverse set of datasets and tasks highlights the general applicability of our approach for unsupervised ranking of segmentation models.
8.1 Electron Microscopy (EM), Mitochondria
@@
MitoEM [17]
@@
The VNC dataset is by far the smallest dataset, therefore, in order to maximise available patches for testing we did not perform a typical test train split. For VNC source models we used the first 18 z-slices for training and the remaining 2 z-slices for validation. We did not transfer VNC based models to a VNC test set, removing this transfer from our set of evaluated transfers. For models trained on alternative sources and transferred to VNC we thus tested on the entire 20 z-slices of VNC.
@@
8.2 Light Microscopy (LM), Nuclei
@@
The SELMA3D 2024 challenge [18].
@@
The set of source models trained on the seven available datasets were then transferred to four target datasets; BBBC039 [38], Hoechst [5], S_BIAD895 [69] and DSB2018 [11] (a fluorescent only subset of the data science bowl 2018 challenge). The target datasets were chosen based on two criteria, firstly, for the quality of their ground-truth annotation, to allow for realistic performance evaluation scores. Secondly, the performance of models transferred to the target dataset should have some differentiation. Model differentiation allows for proper investigation of transferability metrics. If all the models transferred to a dataset perform equally then ranking becomes obsolete.
Instance Segmentation. We trained nuclei instance segmentation models on five datasets; BBBC039 [38], HeLaCytoNuc [14], Hoechst [5], S_BIAD895 [69] and SBIAD1410 [23]. The source models were then transferred to the test sets of four target datasets BBBC039 [38], Hoechst [5], S_BIAD895 [69] and one additional dataset S_BIAD634 [31] was introduced. S_BIAD634 is a dataset of nuclei images of normal and cancer cells from different tissue origins and sample preparation types. The dataset comprises of 42 training images and 37 test images. Once again Target datasets were chosen with the same two criteria for instance ground-truth quality and differentiable model performance under transfer.
8.3 Light Microscopy (LM), Cells

### v3

@@ 5.1 Semantic Segmentation
Tab.˜1 shows all CMR variants exhibit strong linear and monotonic correlations with performance across all target datasets. Comparable results between input- and feature-space perturbations confirm both approaches provide effective ranking. For feature-space perturbations, we examined where in the network they should be applied. Prior work [mi_training-free_2022, he_reshaping_2014] suggests intermediate layers, where representations contain the most redundant information. To test generality, we applied TTD either to all layers (Mito) or only to the bottleneck (Nuclei). As shown in Tab.˜1 and the Supp., both strategies yield similarly strong correlations, but with different effective ranges.
Our model sets include networks trained both with and without data augmentations; the consistently high CMR correlations indicate that ranking is unaffected by source-domain augmentation strategies. “Hard” (CMR-NHD) and “soft” (CMR-EI) metrics perform equivalently well for semantic segmentationlikely because the large number of pixel-wise predictions per image provides a strong signal for the ranking score. UE methods [mi_training-free_2022] typically require multiple inference runs (). In contrast, we find that measuring consistency with a single perturbation per test image is sufficient for reliable model ranking, requiring only two inference passes per image. For both input and feature-space perturbations, a broad range of perturbation strengths produces stable rankings (see Supp.); therefore all main tables report results for a single perturbation setting.
Fig.˜3 illustrates CMR performance on the EPFL [lucchi_learning_2013] mitochondria segmentation dataset across a diverse set of architectures (U-Net, Residual U-Net, UNETR) trained on multiple source datasets with and without input augmentations. Although models perform similarly on their respective source datasets, transfer performance varies substantially, and simple source–target similarity does not reliably predict ranking. For example, UNETR (2) and U-Net (3), trained on the same Rmito dataset, show markedly different target performance. CMR captures both architectural properties and source-domain effects through the learned weights, directly assessing model suitability for the target dataset. While CMR generally correlates strongly with transfer performance, pathological cases can cause mis-rankings. These occur primarily among low-performing transfers where models exhibit “mode collapse” (predicting all foreground or background), producing artificially high consistency. Such cases could be mitigated by combining consistency with cross-model prediction diversity to distinguish stable models from collapsed predictions.
| Class | CMR-EI (Gauss) | CMR-NHD (Gauss) | CMR-EI (DropOut) | CMR-NHD (DropOut) | CMR-EI (Gamma) | CMR-NHD (Gamma) | TS | NuNo | Dispersion | ||||||||||
| IACs | K | 0.90 | (**) | 0.81 | (**) | 0.87 | (*) | 0.87 | (*) | 1.00 | (**) | 0.90 | (*) | 0.21 | (0.2) | 0.14 | (0.4) | 0.07 | (0.4) |
| S | 0.96 | (**) | 0.93 | (*) | 0.94 | (**) | 0.94 | (**) | 1.00 | (**) | 0.96 | (**) | 0.22 | (0.1) | 0.14 | (0.2) | 0.17 | (0.3) | |
| P | 0.99 | (**) | 0.98 | (**) | 0.83 | (*) | 0.80 | (0.1) | 0.99 | (**) | 0.98 | (**) | 0.45 | (0.1) | 0.01 | (0.4) | 0.08 | (0.3) | |
| Teeth | K | 0.81 | (**) | 1.00 | (**) | 0.67 | (0.1) | 0.75 | (0.1) | 0.71 | (*) | 0.90 | (**) | 0.36 | (0.1) | 0.07 | (0.3) | 0.43 | (*) |
| S | 0.89 | (*) | 1.00 | (**) | 0.77 | (0.1) | 0.77 | (0.1) | 0.86 | (*) | 0.96 | (**) | 0.55 | (0.1) | 0.21 | (0.2) | 0.60 | (*) | |
| P | 0.95 | (**) | 0.98 | (**) | 0.85 | (**) | 0.94 | (**) | 0.89 | (**) | 0.95 | (**) | 0.55 | (*) | -0.01 | (0.4) | 0.49 | (0.1) | |
| Mand. | K | 0.62 | (*) | 0.52 | (0.2) | 0.87 | (*) | 0.87 | (**) | 0.43 | (0.3) | 0.81 | (**) | 0.09 | (0.3) | 0.07 | (0.3) | 0.18 | (0.2) |
| S | 0.82 | (*) | 0.68 | (0.1) | 0.87 | (*) | 0.94 | (*) | 0.54 | (0.2) | 0.89 | (**) | 0.17 | (0.2) | 0.12 | (0.3) | 0.25 | (0.2) | |
| P | 0.73 | (0.1) | 0.65 | (0.1) | 0.89 | (**) | 0.96 | (**) | 0.74 | (0.1) | 0.80 | (*) | 0.02 | (0.3) | 0.23 | (0.2) | 0.38 | (0.1) | |
| Sinus. | K | 0.43 | (0.3) | 0.81 | (*) | 0.87 | (*) | 0.83 | (*) | 0.43 | (0.2) | 0.90 | (**) | 0.57 | (*) | -0.07 | (0.3) | 0.34 | (0.1) |
| S | 0.50 | (0.3) | 0.89 | (*) | 0.84 | (*) | 0.94 | (*) | 0.57 | (0.2) | 0.96 | (**) | 0.64 | (*) | -0.05 | (0.4) | 0.35 | (0.2) | |
| P | 0.38 | (0.4) | 0.73 | (0.1) | 0.88 | (**) | 0.90 | (**) | 0.38 | (0.4) | 0.98 | (**) | 0.53 | (0.1) | -0.01 | (0.4) | 0.54 | (*) | |
| Overall | K | 0.71 | (*) | 0.90 | (**) | 0.93 | (*) | 0.73 | (0.1) | 0.81 | (**) | 1.00 | (**) | 0.33 | (0.1) | 0.25 | (0.1) | 0.28 | (0.1) |
| S | 0.86 | (*) | 0.96 | (**) | 0.95 | (*) | 0.83 | (0.1) | 0.89 | (*) | 1.00 | (**) | 0.49 | (0.2) | 0.28 | (0.2) | 0.50 | (0.1) | |
| P | 0.93 | (**) | 0.98 | (**) | 0.93 | (**) | 0.93 | (**) | 0.90 | (**) | 0.96 | (**) | 0.55 | (0.2) | 0.26 | (0.2) | 0.44 | (0.1) | |
Tab.˜2 shows highly multi-class 3D experiments on a public challenge dataset ToothFairy2 [2024TMI, 2025CVPR]. Here, both our CMR-EI and CMR-NHD metrics remain reliable, although CMR-NHD is slightly better, likely due to poor model calibration on the challenge test set [ovadia_can_2019]. Correlation scores are computed after aggregating semantically similar classes, i.e., averaging left/right Inferior Alveolar Canals into IACs, the 32 teeth into Teeth, and left/right Maxillary Sinuses into Sinus. The Overall score shows the mean across all 42 classes in the dataset. Tab.˜2 also reports three competitor metrics adapted to the 3D multi-class setting; since these metrics are inherently multi-class, we report the correlation between the overall metric ranking and the corresponding per-group model performance. Once again, both input and feature space perturbations employed by our method yield the strongest and most coherent rank correlations. Splitting the class-grouped results reveals our metrics are effective and stable across a morphologically heterogeneous set of classes. Our results largely reproduce the challenge rankings without access to target labels, suggesting a practical approach for self-assessing challenge submissions when the test dataset is available. Notably, our metrics are substantially more memory efficient than competitors: also, the CMR-EI variant only requires the maximum probability of each voxel, i.e., predicted class, whereas TS, NuNo, and Dispersion require storing all softmax probabilities, which is prohibitive in 3D (e.g., more than 200GB for ToothFairy2 dataset).
UDA Validation. UDA may be applied to improve model performance, but still requires ranking the adapted models. Tab.˜4 shows ranking correlation scores across four mitochondria datasets and two UDA methods: Mean Teacher (MT) [tarvainen_mean_2018] self-supervised training and adapted batch normalisation (AdaBN) [li_adaptive_2018]. We observe strong correlation between CMR and post-UDA F1-score on EPFL, Hmito and Rmito for both MT and AdaBN, outperforming the previous SOTA for UDA validation, Transfer Score (TS) [yang_can_2024]. VNC exhibits lower correlation scores for both CMR and TS metrics, particularly under the MT approach. For CMR, correlation scores are reduced by a few pathological models, all with EPFL source (see Supp.), that consistently segment non-mitochondrial structures in the target data, yielding stable but incorrect predictions and, consequently, low task performance. Consistency regularisation in MT training reinforces this class confusion between visually similar, but distinct structures present in EPFL and VNC, violating the assumption that source and target models address the same task. Hence, models can be ranked well under direct transfer but exhibit outlier behaviour after MT self-training. This is further highlighted by the AdaBN UDA approach, which lacks a consistency regularisation component and thus remains less affected, resulting in higher correlation scores. Post-UDA ranking is inherently more challenging, as target data becomes part of the model’s adapted ‘source’ distribution, potentially reducing the impact of perturbations used in CMR estimation, yet our approach maintains good correlation.
@@
| Metric | Cells | Nuclei | ||||
| CMR-ARS (Gauss) | K | 0.69 | ±0.15 | 0.72 | ±0.25 | |
| S | 0.83 | ±0.09 | 0.83 | ±0.17 | ||
| P | 0.90 | ±0.04 | 0.79 | ±0.21 | ||
| CMR-ARS (DropOut) | K | 0.69 | ±0.09 | 0.80 | ±0.28 | |
| S | 0.83 | ±0.06 | 0.85 | ±0.24 | ||
| P | 0.91 | ±0.04 | 0.84 | ±0.11 | ||
| SEG | K | 0.12 | ±0.49 | 0.48 | ±0.42 | |
| S | 0.16 | ±0.68 | 0.53 | ±0.49 | ||
| P | 0.31 | ±1.02 | 0.13 | ±0.72 | ||
@@ 5.2 Instance Segmentation
Instance segmentation model ranking remains largely unexplored. Unlike semantic segmentation, instance outputs consist of unordered sets of masks with no fixed correspondence to feature representations, making previously compared transferability metrics inapplicable. Tab.˜4 reports correlation with mean Average Precision (mAP@[0.5:0.95]) [lin_microsoft_2015] for our CMR-ARS method and the ensemble-based SEG method across two tasks and seven target datasets (see Supp. for model details). SEG requires two hyperparameters: the agreement ratio defining ensemble consensus for pseudo ground truth and the centroid dilation radius . Following [sims_seg_2023] we set and estimate as half the mean instance size from a small random sample, reflecting a realistic unsupervised setting. However, SEG proves highly sensitive to these parameters, as reflected by the large variance () values (see Supp.). In contrast, CMR-ARS shows consistently strong correlations with performance under both input and feature-space perturbations, outperforming SEG across all targets. A key advantage of CMR-ARS is that it evaluates instance segmentations directly at the pixel level using the Rand score. In contrast, SEG reduces each instance to a circularly dilated centroid and measures object detections only, which is less sensitive to pixel-level quality differences and problematic for elongated or irregular biomedical objects.
Since CMR relies only on output consistency, it can compare models with arbitrary instantiation strategies. Fig.˜4 illustrates a realistic ranking scenario on the Covid-IF cell segmentation dataset using several widely used models: SAM [archit_segment_2025] with 3 transformer backbones (ViT-T, ViT-B, ViT-L), Cellpose-SAM [pachitariu_cellpose-sam_2025], and the specialist U-Net model “Powerful-Chipmunk” [pape_microscopy-based_2021] from the BioImage Model Zoo [ouyang_bioimage_2022]. Despite substantial differences in architecture and instance generation strategies, CMR accurately reproduces the true mean Segmentation Accuracy (mSA) [everingham_pascal_2010] ranking across this diverse model set (Fig.˜4), demonstrating effective unsupervised model selection in a realistic repository setting.
@@ 6 Limitations and Conclusion
This work introduces a consistency-based model ranking (CMR) method for semantic and instance segmentation that is source-free, unsupervised, and model-agnostic. Our rankings strongly correlate with true target-domain model performance across diverse segmentation tasks and model types, both under direct application and after unsupervised domain adaptation, while requiring only two inference passes on unlabelled target data. The resulting combination of accuracy, efficiency, and conceptual simplicity makes CMR well suited for practical model selection in repository settings with heterogeneous pretrained models.
Like other transferability metrics, CMR assumes alignment between source and target tasks, which may be violated for highly generalist models without explicit task specification (e.g., and Cellpose-SAM). In addition, the current formulation aligns most closely with pixel-wise evaluation metrics such as F1 or mAP; boundary-sensitive metrics could be incorporated through pixel-importance weighting within the consistency calculation. More broadly, we believe that unsupervised ranking methods such as CMR can enable scalable model reuse and benchmarking in emerging model repositories, supporting wider adoption of segmentation models in domains where labeled data remains scarce.
@@
Appendix 0.A Datasets
The general experimental setup for source model ranking under transfer to a target dataset is as follows, a set of ‘source’ models are trained for the same task (i.e nuclei semantic segmentation). The models can have a range of architectures and source datasets. A single dataset is then selected as ‘target’ and all of the models from are applied to the test set of the target data. We then rank the models and compare this ranking to the true transfer performance ranking, obtained using ground-truth. In this section we further detail the employed datasets, and in section Appendix˜0.B the tested models.
We experimented applying our consistency based model ranking (CMR) method across a wide range of datasets and modalities from the microscopy and biomedical domains. We focused on ranking the target performance of semantic and instance segmentation models across several distinct tasks. The tasks covered arguably the most common problems in microscopy image analysis; nuclei, cell and mitochondria segmentation. We additionally investigated a popular biomedical segmentation challenge (i.e, ToothFairy2) for human jaw segmentation. Our wide range of experiments across such a diverse set of datasets and tasks highlights the general applicability of our approach for unsupervised ranking of segmentation models.
0.A.1 Electron Microscopy (EM), Mitochondria
@@
Hmito, Rmito [franco-barranco_current_2023]
@@
The VNC dataset is by far the smallest dataset, therefore, in order to maximise available patches for testing we did not perform a typical test train split. For VNC source models we used the first 18 z-slices for training and the remaining 2 z-slices for validation. We did not transfer VNC source trained models to a VNC test set, removing this transfer from our set of evaluated transfers. For models trained on alternative sources and transferred to VNC we thus tested on the entire 20 z-slices of VNC.
@@
0.A.2 Light Microscopy (LM), Nuclei
@@
The SELMA3D 2024 challenge [leal-taixe_benchmark_2019].
@@
The set of source models trained on the seven available datasets were then transferred to four target datasets; BBBC039 [ljosa_annotated_2012], Hoechst [arvidsson_annotated_2023], S_BIAD895 [von_chamier_democratising_2021] and DSB2018 [caicedo_nucleus_2019] (a fluorescent only subset of the data science bowl 2018 challenge). The target datasets were chosen based on two criteria, firstly, for the quality of their ground-truth annotation, to allow for realistic performance evaluation scores. Secondly, the performance of models transferred to the target dataset should have some differentiation. Model differentiation allows for proper investigation of ranking metrics. If all the models transferred to a dataset perform equally then ranking becomes obsolete and hard to evaluate.
Instance Segmentation. We trained nuclei instance segmentation models on five datasets; BBBC039 [ljosa_annotated_2012], HeLaCytoNuc [de_helacytonuc_2024], Hoechst [arvidsson_annotated_2023], S_BIAD895/SB-895 [von_chamier_democratising_2021] and SBIAD1410 [hawkins_rescu-nets_2024]. The source models were then transferred to the test sets of four target datasets BBBC039 [ljosa_annotated_2012], Hoechst [arvidsson_annotated_2023], S_BIAD895/SB-895 [von_chamier_democratising_2021] and one additional dataset S_BIAD634/SB-634 [kromp_annotated_2020] was introduced. S_BIAD634/SB-634 is a dataset of nuclei images of normal and cancer cells from different tissue origins and sample preparation types. The dataset comprises of 42 training images and 37 test images. Once again Target datasets were chosen with the same two criteria for instance ground-truth quality and differentiable model performance under transfer. For details on training of instance segmentation models see Appendix˜0.B.
0.A.3 Light Microscopy (LM), Cells